# Identifying Astrophysical Anomalies in 99.6 Million Source Cutouts from the Hubble Legacy Archive Using AnomalyMatch

### v1

@@ Abstract
Astronomical archives contain vast quantities of unexplored data that potentially harbour rare and scientifically valuable cosmic phenomena. We have systematically searched approximately 100 million image cutouts from the entire Hubble Legacy Archive using the recently developed AnomalyMatch method, which combines semi-supervised and active learning techniques for the efficient detection of astrophysical anomalies. This comprehensive search rapidly uncovered a multitude of astrophysical anomalies presented here that significantly expand the inventory of known rare objects. Among our discoveries are 138 new candidate gravitational lenses, 18 jellyfish galaxies, and 417 mergers or interacting galaxies. The efficiency and accuracy of our iterative detection strategy allows us to trawl the complete archive within just 2–3 days, highlighting its potential for large-scale astronomical surveys. We present a detailed overview of these newly identified objects, discuss their astrophysical significance, and demonstrate the considerable potential of AnomalyMatch to efficiently explore extensive astronomical datasets, including, e.g., upcoming Euclid data releases.
Computational methods (1965) – Surveys (1671) – Galaxies (573) – Amorphous irregular galaxies (37) – Interacting galaxies (802) – Gravitational lensing (670)
@@ 1 Introduction
With the launch of Euclid (Laureijs et al., 2011), the beginning of operations of the Vera C. Rubin observatory (Ivezić et al., 2019) and the construction of the Nancy Grace Roman Space Telescope (Mosby et al., 2020), the size of astronomical datasets is growing rapidly. These observatories will survey a large fraction of the sky in an agnostic way, leading to the potential discovery of a large number of new objects of astrophysical interest allowing us to expand our catalogues of rare galaxy morphology types, including cosmological effects like gravitational lenses or galaxies undergoing the effects of dense environments.
@@
What many of these rare systems have in common are their methods of discovery. Often they are found by experts ‘manually’ exploring their data, and detecting odd morphologies or serendipitously finding objects of interest (Pearce-Casey et al., 2024; Acevedo Barroso et al., 2024). Another common approach leverages the contribution of citizen scientists. Using platforms such as Zooniverse from the Galaxy Zoo collaboration (Lintott et al., 2008), volunteers are able to mark any galaxies with anomalous morphologies as ‘odd’.
@@
Approaches such as Astronomaly (Lochner & Bassett, 2021) aim to resolve this problem. Rather than training to find individual types of anomalies, they utilise a combination of isolation forests (Liu et al., 2008) or local outlier factor (Breunig et al., 2000) with active learning to identify anomalies specifically sought by a user. In further study, this can be combined with feature spaces produced by machine learning models (e.g Zoobot; Walmsley et al., 2022, 2023) to identify further examples of an anomaly. Within these feature spaces, objects of similar morphologies appear close together and can be extracted from areas of interest.
In this work we utilise a different approach. Rather than relying on feature spaces, we have developed a neural network-based approach to apply directly to image data of sources. We focus on discovering ‘similar’ anomalies in a binary classification task rather than initially starting from an unsupervised approach. Our method utilises semi-supervised learning (SSL) techniques such as pseudo-labelling (Lee, 2013) and consistency regularization (Bachman et al., 2014) to leverage the limited labelled data available combined with a large number of unlabelled samples to make binary classifications on an anomalous or nominal source.
@@
Our approach, named AnomalyMatch, is fully described in the companion paper (Gomez & O’Ryan, 2025). There, we conduct thorough benchmarking and validation of the approach on established datasets such as mini-ImageNet (Vinyals et al., 2016) and Galaxy MNIST (Walmsley et al., 2022). In parallel to the development, we actively applied this approach to searching for anomalies in the entire Hubble Space Telescope Legacy Archive [^1] HLA: https://hla.stsci.edu/ (HLA).
@@
We use the European Space Agency’s (ESA) science data platform ESA Datalabs [^2] ESA Datalabs: https://datalabs.esa.int/ (Navarro et al., 2024) to conduct this search. In initial testing, we attempted to identify edge-on protoplanetary disks. These systems form very distinct ‘hamburger’ or ‘butterfly’ shapes which can easily be detected by image classification algorithms. However, only very few of them are known in the literature. As we developed and updated the algorithm, we expanded our search via active learning and marking other objects of astrophysical interest. In this work, we discuss and display the wealth of new objects found in the development and application of AnomalyMatch and discuss the implications for our model. Overall, we report 1,176 newly found anomalies which span 19 different classes. These include galaxy mergers, gravitational lenses and arcs, edge-on protoplanetary disks, and a host of rare galaxy morphologies.
@@
Hereafter, an astrophysical anomaly (or simply anomaly) is any galaxy or object which shows characteristics of an outlier from the general population. This primarily involves a difference in morphology that we are able to detect.
@@ 2.1 Source Cutouts of the HLA
For each cutout, a fixed size of 150$\times$150 pixels was used (7.5”$\times$7.5”), with a LinearStretch and ZScaleInterval from Astropy (Astropy Collaboration et al., 2013, 2018, 2022). Each image was stored as a one-channel greyscale JPEG image. While this does lead to a loss of quality of the image, it allows for more efficient storage. The image creation process was prioritised such that the morphological features of each galaxy were visible. This approach, however, was not optimised to find low surface brightness features, potentially impacting our detection of anomalies such as jellyfish galaxies.
@@ 2.2 Training Set
Our initial goal was to find more examples of edge-on protoplanetary disks. Therefore, we started our search with three examples of these anomalies, a set of labelled nominal data and $\approx$99.6 million unlabelled images. Nominal imaging was selected by visual inspection of randomly selected images from full dataset. We selected images of individual galaxies, star fields and any imaging artifacts that appeared in the sample. Figure 1 shows the three anomalies we initially used to train AnomalyMatch. However, upon conducting active learning, we found that the odd morphology of these systems led to the discovery of other objects of interest also being given high anomaly scores.
@@
We began to build a larger and general training set of different anomalies as we continued active learning. The anomalies at this stage were classified based on their morphology alone, rather than by searching the literature for specific kinds of objects. In total, the training set consisted of 1400 images: 375 anomalies and 1025 nominal images. The anomalies were primarily mergers (178); galaxies which are undergoing close interactions and coalescing. However, we also uncovered many gravitational lenses that were incorporated into the training set (63). Figure 2 shows a subset of the training set that was built during the development of AnomalyMatch.
@@ 3.1 AnomalyMatch
AnomalyMatch is an approach combining active learning and SSL that we have developed and optimised for the purpose of identifying anomalies in large datasets. It is fully described, benchmarked, and validated in the companion paper Gomez & O’Ryan (2025). We provide a brief description of it here.
@@
AnomalyMatch improved on MSMatch by incorporating an active learning loop. The user provides AnomalyMatch with a set of labelled and unlabelled data, and then conducts an initial round of training. Predictions are then made on the unlabelled data, and an ‘anomaly score’ is given to each image based on the model’s classification. The images are ranked by this score, and then shown to the user in an easy to use UI. An example of this is shown in Gomez & O’Ryan (2025). The user can then provide additional labels for consecutive training on a dataset that is progressively expanded and refined. This process of active learning is where the the bulk of our initial training set has been collected. Figure 3 shows the full AnomalyMatch workflow. The AnomalyMatch code will be published on GitHub pending completion of an ESA open-source licensing process. [^3] AnomalyMatch: https://github.com/esa/AnomalyMatch.
@@ 3.2 Anomaly Identification
To identify anomalies, the only initial information we have is their morphology. During active learning and inference over the HLA, there was no direct access to further ancillary data. The underlying distribution of anomaly score is highly bimodal - the model is very confident that an object is anomalous or not. Given the strong class imbalance and training on only a limited subset of data, model scores were not calibrated and they are not probabilities. We do not apply a cut to this score but investigate their ranked order.
After training on all labelled data and a subset of the unlabelled data, we visually inspect the 5000 sources with the highest predicted anomaly scores of the entire HLA. Based on either the morphology of the galaxy or the object in the image, we are able to make an initial estimation of the kind of anomaly we find. However, to confirm or reject this, we must turn to the literature. We made use primarily of SIMBAD [^4] https://simbad.u-strasbg.fr/simbad/ and the ESASky platform [^5] https://sky.esa.int/esasky/. SIMBAD is an excellent tool to programmatically check if the anomalies we found have any associated papers. We conducted a cone search about the coordinates of each anomaly with a radius of $3"$. We then checked if any source within this radius had was associated to any work at the time of writing. This could then be followed up quickly and efficiently using ESASky.
@@ 3.2.3 Gravitational Lenses
The most recent works searching for gravitational lenses used Euclid data and have employed both expert labelling (Pearce-Casey et al., 2024) and machine learning techniques (Euclid Collaboration et al., 2025a). We classify a gravitational lens about an object when we observe these structures around the galaxies in our sample.
@@ 3.2.7 Clumpy Galaxies
Clumpy galaxies, owing their name to the presence of large and luminous star-forming clumps within them, are gas rich systems which often reside at $31<z<$ 3 (Bournaud et al., 2008; Shibuya et al., 2016). Samples typically range from 100 - 1000 systems (Guo et al., 2015; Adams et al., 2022). The clumps can easily be mistaken for a second core within the galactic disk leading to possible misclassification as a merger. We ensure that when making either the merger or clumpy classification, we consider for any morphological disturbance to the galactic disk. Also, the existence of more than one secondary core in the galaxy is likely an indication that it is a clumpy galaxy rather than a merger.
@@ 3.2.8 Active Galactic Nuclei
3.2.9 Ring & Collisional Ring Galaxies
@@ 3.2.12 High Redshift Galaxies
3.2.13 Odd Galaxies & Unknown Galaxies
@@
3.3 Anomaly Score & SHAP Distributions
@@
Figure 5 shows the resultant distribution of anomaly scores across all five HDF5 files ($\approx 500,000$ sources). The majority of our sources are given scores $\approx 0$. We do not expect many anomalies to appear in a random selection of $0.5\%$ of the data. At an anomaly score of $\approx 1$, we observe anomalies which could be of interest.
@@ 4 Results
To perform the anomaly sub-classification, we initially selected the top 5000 scored anomalies from the HLA. First, we apply de-duplication to these 5000 samples. The output from AnomalyMatch is a CSV file containing the filename of the source and the score given to the source. In our case, the source filename is the source ID found in the HSC. We cross match each of our sources with the HSC based on this, and extract their coordinates. We then apply an aggressive radial cross match between each of the sources within $10"$.
We apply such a large cut as the likelihood that two of our anomalies are within this separation is low. It allows us to completely de-duplicate the final sample, and ensure we are not using expert time re-inspecting duplicate entries. This reduces the number of anomalies from 5,000 to 1,339 unique images. This also highlights the high level of shredding and duplication in the HSC. Upon de-duplication, DOR acted as the expert classifier and classified each anomaly according to the system defined in Section 3.
@@ 4.1 Detected Anomalies
The unique 1,339 anomalies taken from the ranking of highest anomaly score were classified into their sub-classifications. This was done by a combination of accounting for morphology and searching the literature for works related to each anomaly. Table 1 shows a breakdown of these classification of the anomalies. Figure 7 shows a representative sample of all sub-classifications where we found at least five anomalies, except for our ‘odd’, ‘nominal’ and ‘unknown’ classifications, which will be presented later in the section. The ‘nominal’ classification is our false positive rate, where AnomalyMatch returned a normal source.
@@
We do not include our sample of 39 gravitational arcs from our lens classification here. Often, these arcs - formed by clusters rather than galaxies - are so large, that they extend out of the field of view of our source cutouts. We see the true extent of these systems in ESASky, although only parts of them are flagged as anomalies here. Other cutouts containing the arc are removed in our aggressive cross matching and de-duplication of $10"$. We also ensure that each of our 39 arcs is unique using ESASky.
@@
We find 35 different jellyfish galaxies in our 1,339 unique systems. These have been classified as such as they clearly have stripping occurring, and they are residing in a dense cluster environment. On each of the examples shown in Figure 7, the bow shock is present in the direction of travel.
@@
AnomalyMatch excelled in detecting anomalies for which we had provided no examples. This allowed us to expand the training set and identify more of these objects. This was particularly true of our lensed quasars. Figure 8 shows the four systems that we found in HLA. We initially provided AnomalyMatch with no examples of these, however, even in the first training iterations, it identified these anomalies and scored them highly. Therefore, we added one to the training set and identified four additional lensed quasars.
@@
Galaxies we classify as nominal by visual inspection, but that received a high score by AnomalyMatch, represent the contamination in our final anomaly sample. We find a contamination rate of $\approx 6\%$. Figure 11 shows a representative example of ten sources. Primarily, these sources are very small sources which exhibit some potential merging behaviour, but it is difficult to make a specific classification. Other examples include sources with artifacts from the observation itself, aligned with a point source and star fields. Star fields with this morphology are often either from the Panchromatic Hubble Andromeda Treasury (Dalcanton et al., 2012) or studies of large globular clusters with the ACS instrument. Therefore, many of the nominal images we find here are anomalous when compared to the rest of the data, but they are of limited value to the astrophysical community.
@@
Table 2: Table showing the first five rows of the catalogue released with this work.
@@
For the ease of use of our catalogue, we also release all of the sources and table on Zenodo (DOI: 10.5281/zenodo.1529864210.5281/zenodo.15298642). Here, the images are contained in separate .zip files based on their anomaly classification. The filenames are the Source IDs of the sources.

### v3

@@ Abstract
Aims. Astronomical archives contain vast quantities of unexplored data that potentially harbour rare and scientifically valuable cosmic phenomena. We leverage new semi-supervised methods to extract such objects from the Hubble Legacy Archive.
Methods. We have systematically searched approximately 100 million image cutouts from the entire Hubble Legacy Archive using the recently developed AnomalyMatch method, which combines semi-supervised and active learning techniques for the efficient detection of astrophysical anomalies. This comprehensive search rapidly uncovered a multitude of astrophysical anomalies presented here that significantly expand the inventory of known rare objects.
Results. Among our discoveries are 86 new candidate gravitational lenses, 18 jellyfish galaxies, and 417 mergers or interacting galaxies. The efficiency and accuracy of our iterative detection strategy allows us to trawl the complete archive within just 2–3 days, highlighting its potential for large-scale astronomical surveys.
Conclusions. We present a detailed overview of these newly identified objects, discuss their astrophysical significance, and demonstrate the considerable potential of AnomalyMatch to efficiently explore extensive astronomical datasets, including, e.g., upcoming Euclid data releases.
Key Words.:
Surveys – Galaxies – Galaxies: irregular – Galaxies: interacting – Gravitational lensing: strong
@@ 1 Introduction
With the launch of Euclid (Euclid Collaboration et al., 2025b), the beginning of operations of the Vera C. Rubin observatory (Ivezić et al., 2019) and the construction of the Nancy Grace Roman Space Telescope (Mosby et al., 2020), the size of astronomical datasets is growing rapidly. These observatories will survey a large fraction of the sky in an agnostic way, leading to the potential discovery of a large number of new objects of astrophysical interest allowing us to expand our catalogues of rare galaxy morphology types, including cosmological effects like gravitational lenses or galaxies undergoing the effects of dense environments.
@@
What many of these rare systems have in common are their methods of discovery. Often they are found by experts ‘manually’ exploring their data, and detecting odd morphologies or serendipitously finding objects of interest (Pearce-Casey et al., 2025; Acevedo Barroso et al., 2025). Another common approach leverages the contribution of citizen scientists. Using platforms such as Zooniverse from the Galaxy Zoo collaboration (Lintott et al., 2008), volunteers are able to mark any galaxies with anomalous morphologies as ‘odd’.
@@
Approaches such as Astronomaly (Lochner & Bassett, 2021) aim to resolve this problem. Rather than training to find individual types of anomalies, they utilise a combination of isolation forests (Liu et al., 2008) or local outlier factor (Breunig et al., 2000) with active learning to identify anomalies specifically sought by a user. In further study, this can be combined with feature spaces produced by machine learning models (e.g Zoobot; Walmsley et al., 2022, 2023) to identify further examples of an anomaly. This is effectively done with the ulisse tool (Doorenbos et al., 2022), where these feature spaces can be analysed such that objects of similar morphologies appear close together and can be extracted from areas of interest.
Related works in the area often use unsupervised learning to search for astrophysical anomalies. Unsupervised methods in this context leverage unlabelled data and its data structure to create groups of similar objects expecting anomalies to cluster or fall outside clusters. Such an approach was successfully applied in Stein et al. (2022). In this work, a self-supervised machine learning method was used to mine the entire Dark Energy Spectroscopic Instrument survey Data Release 9 for strong gravitational lenses. Their self-supervised method created an output representation space which was linked to the morphology of the objects trained upon. Once this feature space was created, the representation vector of a single image of a gravitational lens (output by the same model) was then used to conduct a similarity search through 76 million images, identifying 1,192 candidate gravitational lenses with excellent computational efficiency.
In this work we utilise a different approach. We frame anomaly detection as an imbalanced binary classification problem. This is not a fundamentally different task as anomaly detection inherently involves distinguishing rare instances from common ones, but it allows us to leverage semi-supervised learning techniques specifically designed for extreme class imbalance, enabling effective learning from minimal labelled anomaly examples. Our method utilises semi-supervised learning (SSL) techniques such as pseudo-labelling (Lee, 2013) and consistency regularization (Bachman et al., 2014) to make binary classifications on an anomalous or nominal source. Additionally, we use active learning, where experts iteratively validate and label predictions of the top ranking scored images. This hybrid approach offers key advantages: it requires very few initial labels (even fewer than ten anomalies), enables expert validation throughout the process, and effectively utilises both unlabelled data and expert knowledge to progressively improve detection performance.
@@
Our approach, named AnomalyMatch, is fully described in the companion paper (Gómez & O’Ryan, 2025). There, we conduct thorough benchmarking and validation of the approach on established datasets such as mini-ImageNet (Vinyals et al., 2016) and Galaxy MNIST (Walmsley et al., 2022). In parallel to the development, we actively applied this approach to searching for anomalies in the entire Hubble Space Telescope Legacy Archive [^1] HLA: https://hla.stsci.edu/ (HLA).
@@
We use the European Space Agency’s (ESA) science data platform ESA Datalabs [^2] ESA Datalabs: https://datalabs.esa.int/ (Navarro et al., 2024) to conduct this search. In initial testing, we attempted to identify edge-on protoplanetary disks. These systems form very distinct ‘hamburger’ or ‘butterfly’ shapes which can easily be detected by image classification algorithms. However, only very few of them are known in the literature. As we developed and updated the algorithm, we expanded our search via active learning and marking other objects of astrophysical interest. This was due to multiple factors. First, was a limitation of the data we are using. We search primarily in a previously created dataset of $F814W$ filter HST sources observed with the Advanced Camera for Surveys/Wide Field Channel (ACS/WFC; hereafter only ACS). In this dataset, the number of edge-on protoplanetary disks is limited as they are often observed at other wavelengths, in the Wide-Field Planetary Camera-2 (WFPC2) or are very difficult to resolve in HST data (Burrows et al., 1996; Krist et al., 1998; Stapelfeldt et al., 2003; Ricci et al., 2008; Angelo et al., 2023, e.g. see). The second was that our algorithm began to detect numerous other objects of interest that we wished to pursue. These included lensed quasars, mergers and gravitational lenses.
In this work, we discuss and display the wealth of new objects found in the development and application of AnomalyMatch and discuss the implications for our model. Overall, we report 1,176 newly found anomalies which span 19 different classes. These include galaxy mergers, gravitational lenses and arcs, edge-on protoplanetary disks, and a host of rare galaxy morphologies.
@@
For an excellent description of an anomaly, see the introduction of Ruff et al. (2020). In this work, we define an astrophysical anomaly (or simply anomaly) to be an astrophysical object that shows morphological characteristics which deviate notably from the general population.
@@ 2.1 Source cutouts of the HLA
For each cutout, a fixed size of $\times150$ pixels was used ($\times7.5$”), with a LinearStretch and ZScaleInterval from Astropy (Astropy Collaboration et al., 2013, 2018, 2022). Each image was stored as a one-channel greyscale JPEG image. While this does lead to a loss of quality of the image, it allows for more efficient storage. The image creation process was prioritised such that the morphological features of each galaxy were visible. This approach, however, was not optimised to find low surface brightness features, potentially impacting our detection of anomalies such as jellyfish galaxies.
@@ 2.2 Training set
Our initial goal was to find more examples of edge-on protoplanetary disks. Therefore, we started our search with three examples of these anomalies, a set 128 labelled nominal data and $\approx$99.6 million unlabelled images. The nominal imaging was selected by visual inspection of randomly selected images from the full dataset. We selected images of individual galaxies, star fields and any imaging artifacts that appeared in the sample. Figure 1 shows the three anomalies we initially used to train AnomalyMatch. However, upon conducting active learning, we found that the odd morphology of these systems led to the discovery of other objects of interest also being given high anomaly scores. With these new objects, we began to pivot our anomaly search to other interesting objects which were also more likely to appear in $F814W$ data.
@@
From this attempt to search for edge-on protoplanetary disks, we had constructed a second catalogue of interesting objects through active learning. The anomalies at this stage were classified based on their morphology alone, rather than by searching the literature for specific kinds of objects. In total, the training set consisted of 1400 images: 375 anomalies and 1025 nominal images. We continued building the nominal labelled set during active learning, adding to it substantially to represent further un-interesting morphologies and unaccounted artifacts that were scored highly. The anomalies were primarily mergers (178); galaxies which are undergoing close interactions and coalescing. However, we also uncovered many gravitational lenses that were incorporated into the training set (63). Figure 2 shows a subset of the training set that was built during the development of AnomalyMatch.
@@ 3.1 AnomalyMatch
AnomalyMatch is an approach combining active learning and SSL that we have developed and optimised for the purpose of identifying anomalies in large datasets. It is fully described, benchmarked, and validated in the companion paper Gómez & O’Ryan (2025). We provide a brief description of it here.
@@
AnomalyMatch improved on MSMatch by incorporating an active learning loop. The user provides AnomalyMatch with a set of labelled and unlabelled data, and then conducts an initial round of training. Predictions are then made on the unlabelled data, and an ‘anomaly score’ is given to each image based on the model’s classification. The images are ranked by this score, and then shown to the user in an easy to use UI. An example of this is shown in Gómez & O’Ryan (2025). The user can then provide additional labels for consecutive training on a dataset that is progressively expanded and refined. This process of active learning is where the the bulk of our initial training set has been collected. Figure 3 shows the full AnomalyMatch workflow. The AnomalyMatch code will be published on GitHub pending completion of an ESA open-source licensing process. [^3] AnomalyMatch: https://github.com/esa/AnomalyMatch.
@@ 3.2 Anomaly identification
To identify anomalies, the only initial information we have is their morphology. During active learning and inference over the HLA, there was no direct access to further ancillary data. The underlying distribution of anomaly score is highly skewed to nominal images - meaning the model is very confident that an object is not anomalous. Given the strong class imbalance and training on only a limited subset of data, model scores were not calibrated and they are not probabilities. We do not apply a cut to this score but investigate their ranked order.
After training on all labelled data and a subset of the unlabelled data, we visually inspect the 5000 sources with the highest predicted anomaly scores of the entire HLA. Based on either the morphology of the galaxy or the object in the image, we are able to make an initial estimation of the kind of anomaly we find. However, to confirm or reject this, we must turn to the literature. We made use primarily of SIMBAD [^4] https://simbad.u-strasbg.fr/simbad/ and the ESASky platform [^5] https://sky.esa.int/esasky/. SIMBAD is an excellent tool to programmatically check if the anomalies we found have any associated papers. We conducted a cone search about the coordinates of each anomaly with a radius of 33". We then checked if any source within this radius had was associated to any work at the time of writing. This could then be followed up quickly and efficiently using ESASky.
@@ 3.2.3 Gravitational lenses
The most recent works searching for gravitational lenses used Euclid data and have employed both expert labelling (Pearce-Casey et al., 2025) and machine learning techniques (Euclid Collaboration et al., 2025a). We classify a gravitational lens about an object when we observe these structures around the galaxies in our sample.
@@ 3.2.7 Clumpy galaxies
Clumpy galaxies, owing their name to the presence of large and luminous star-forming clumps within them, are gas rich systems which often reside at $1<z<3$ (Bournaud et al., 2008; Shibuya et al., 2016). Samples typically range from 100 - 1000 systems (Guo et al., 2015; Adams et al., 2022). The clumps can easily be mistaken for a second core within the galactic disk leading to possible misclassification as a merger. We ensure that when making either the merger or clumpy classification, we consider for any morphological disturbance to the galactic disk. Also, the existence of more than one secondary core in the galaxy is likely an indication that it is a clumpy galaxy rather than a merger.
@@ 3.2.8 Active galactic nuclei
3.2.9 Ring and collisional ring galaxies
@@ 3.2.12 High redshift galaxies
3.2.13 Odd galaxies and unknown galaxies
@@
3.3 Anomaly score and SHAP distributions
@@
Figure 5 shows the resultant distribution of anomaly scores across all five HDF5 files ($\approx 500,000$ sources). The majority of our sources are given scores $\approx 0$. While we do not know the underlying distribution or occurrence rates of anomalies in the dataset, we do not expect many anomalies to appear in a random selection of $0.5\%$ of the data. At an anomaly score of $\approx 1$, we observe anomalies which could be of interest.
@@ 4 Results
To perform the anomaly sub-classification, we initially selected the top 5000 scored anomalies from the HLA. First, we apply de-duplication to these 5000 samples. The output from AnomalyMatch is a CSV file containing the filename of the source and the score given to the source. In our case, the source filename is the source ID found in the HSC. We cross match each of our sources with the HSC based on this, and extract their coordinates. We then apply an aggressive radial cross match between each of the sources within 1010".
We apply such a large cut as the likelihood that two of our anomalies are within this separation is low. It allows us to completely de-duplicate the final sample, and ensure we are not using expert time re-inspecting duplicate entries. This reduces the number of anomalies from 5,000 to 1,338 unique images. This also highlights the high level of shredding and duplication in the HSC. Upon de-duplication, DOR acted as the expert classifier and classified each anomaly according to the system defined in Section 3.
@@ 4.1 Detected anomalies
The unique 1,338 anomalies taken from the ranking of highest anomaly score were classified into their sub-classifications. This was done by a combination of accounting for morphology and searching the literature for works related to each anomaly. Table 1 shows a breakdown of these classification of the anomalies. Figure 7 shows a representative sample of all sub-classifications where we found at least five anomalies, except for our ‘odd’, ‘nominal’ and ‘unknown’ classifications, which will be presented later in the section. The ‘nominal’ classification is our false positive rate, where AnomalyMatch returned a normal source.
@@
We do not include our sample of 39 gravitational arcs from our lens classification here. Often, these arcs - formed by clusters rather than galaxies - are so large, that they extend out of the field of view of our source cutouts. We see the true extent of these systems in ESASky, although only parts of them are flagged as anomalies here. Other cutouts containing the arc are removed in our aggressive cross matching and de-duplication of 1010". We also ensure that each of our 39 arcs is unique using ESASky.
@@
We find 37 different jellyfish galaxies in our 1,338 unique systems. These have been classified as such as they clearly have stripping occurring, and they are residing in a dense cluster environment. On each of the examples shown in Figure 7, the bow shock is present in the direction of travel.
@@
AnomalyMatch successfully detected several types of anomalies for which we had provided no explicit training examples, such as lensed quasars. While these unseen anomaly types likely share morphological features with our trained classes (e.g., lensed quasars may exhibit arc-like features similar to gravitational lenses or multiple bright regions like mergers), their detection demonstrates the method’s ability to generalise to morphologically similar but distinct anomaly types. However, we acknowledge that unseen anomalies with fundamentally different morphologies from our training set would likely be missed. Figure 8 shows the four systems that we found in HLA. We initially provided AnomalyMatch with no examples of these, however, even in the first training iterations, it identified these anomalies and scored them highly. Therefore, we added one to the training set and identified four additional lensed quasars.
@@
Galaxies we classify as nominal by visual inspection, but that received a high score by AnomalyMatch, represent the contamination in our final anomaly sample. We find a contamination rate of $\approx 10\%$. Figure 11 shows a representative example of ten sources. Primarily, these sources are very small sources which exhibit some potential merging behaviour, but it is difficult to make a specific classification. Other examples include sources with artifacts from the observation itself, aligned with a point source and star fields. Star fields with this morphology are often either from the Panchromatic Hubble Andromeda Treasury (Dalcanton et al., 2012) or studies of large globular clusters with the ACS instrument. Therefore, many of the nominal images we find here are anomalous when compared to the rest of the data, but they are of limited value to the astrophysical community.
@@
Table 2: First five rows of the MRT released with this work. [^6] The table contains the (1) Source ID from the HSC, (2) Right Ascension in degrees, (3) Declination in Degrees, (4) classification in this work and (5) whether the classification was made using the literature.
@@
For the ease of use of our catalogue, we also release all of the sources and table on Zenodo (DOI: 10.5281/zenodo.15298641). Here, the images are contained in separate .zip files based on their anomaly classification. The filenames are the Source IDs of the sources.